OpenStudy (anonymous):

Which of the following inequalities matches the graph? -6x + y < 3 6x + y < 3 6x - y < -3 The correct inequality is not listed

OpenStudy (ybarrap):

Posting the graph?

OpenStudy (anonymous):

im trying

OpenStudy (ybarrap):

ok

OpenStudy (anonymous):

i know is hard to read but can someone please walk me through it

HanAkoSolo (jamierox4ev3r):

also, it may help to turn all of these into slope intercept form... y=mx+b or in this case, y< or >mx+b

HanAkoSolo (jamierox4ev3r):

can you convert the given equations to this form @boycrazzy ?

OpenStudy (ybarrap):

let's find two points on the line, looks like y(0)=3 and y(1)=8, can you give us some other points to confirm?

OpenStudy (anonymous):

no can you show me

HanAkoSolo (jamierox4ev3r):

for example -6x+y<3 +6x +6x ---------- y<6x+3 That is for the first equation

OpenStudy (anonymous):

which points @ybarrap

OpenStudy (ybarrap):

y(0) = b in @Jamierox4ev3r 's equation, btw. Now you just need slope (m)

OpenStudy (ybarrap):

On your graph, find an x and corresponding y value.

OpenStudy (anonymous):

oh 3,4

HanAkoSolo (jamierox4ev3r):

if you look at this, 6 is the slope, and 3 is the y-intercept (where the slope intersects the y-axis. Be sure to note that for this particular equation, both of the values are positive

HanAkoSolo (jamierox4ev3r):

@ybarrap and yes that is right

OpenStudy (anonymous):

Thanks you i think i know the answer now is it -6x + y < 3

OpenStudy (ybarrap):

That looks right.

OpenStudy (anonymous):

Thanks alot can you also help me with a similar question

OpenStudy (ybarrap):

Hopefuly the graph is easier to see \( \ddot\smile\)

OpenStudy (anonymous):

Which of the following inequalities matches the graph? The correct inequality is not listed. 5x + y 1 5x + y 1 5x - y 1

OpenStudy (anonymous):

5x + y >= 1 5x + y <= 1 5x - y >= 1

OpenStudy (ybarrap):

Pick out a couple of easy points: y(1)=4 and y(0) = -1 look good.

OpenStudy (ybarrap):

Next find the slope: m = (4 - (-1) ) / (1 - 0) = 5

OpenStudy (ybarrap):

now we have y = 5x + b. We can find b by plugging in x=1 and y=4, for instance, and solve for b. Want to try that?

OpenStudy (ybarrap):

I get b = -1, is that what you get?

OpenStudy (ybarrap):

We now have the line that divides the shaded region from the unshaded

OpenStudy (anonymous):

i dont really get much of what your saying cause i saw this equations a different way than you do im going to show you make work and can you just check it for me

OpenStudy (anonymous):

i solve***

OpenStudy (ybarrap):

y = 5x -1 is our equation. Now is the shaded region smaller or larger than values in this line? For example, take point x=2 and y=1. Is y=1 smaller than a point on the line we just created. That is y = 5(2)-1 = 9. This point on the line is greater than the random point we selected, 1.

OpenStudy (ybarrap):

show me what you got

OpenStudy (ybarrap):

The points in the shaded region all are smaller than points on the line. Do you see that?

OpenStudy (ybarrap):

So y < 5x -1

OpenStudy (ybarrap):

Can you rearrange this to put in the form you need?

OpenStudy (anonymous):

o i just graph that so it would be negative 1 the point on y and than 5 rise and 1 run

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

so is the correct answer 5x + y >= 1

OpenStudy (ybarrap):

Let's work on the inequality. y < 5x -1 y -5x < -1 5x - y > 1 This is our final answer.

OpenStudy (anonymous):

ok i understand thats for helping

OpenStudy (ybarrap):

yw

OpenStudy (anonymous):

haha i just turned in my test scores and the first question you helped me with wasnt right im going to try and figure it out now i still got a 90% anyways good night and thats for the help